# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

### v1

@@ A B $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
Table 21: Accuracy on different weight alignment in PQ-LoRA under Heterogeneous PFL-Static setup on DRAKE. ‘Random’ initializes matrices randomly, while Aligned (MSE) and Aligned (CCA) use matrices aligned by MSE loss and canonical correlation analysis (CCA), respectively.
A.16 Effect of $N_{B}$ in PQ-LoRA
We ablate the effect of $N_{B}$, the number of layers employing PQ-LoRA, and summarize the results in Tab.22. In a model $W$ with $|W|$ layers, increasing $N_{B}$ reduces the layers using conventional LoRA to $|W|-N_{B}$. Since PQ-LoRA has fewer trainable parameters (i.e., $r^{2}+r$) than conventional LoRA (i.e., $r\times(d_{I}+d_{O})$), increasing $N_{B}$ decreases the total number of learnable parameters in $W$, as $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$ are frozen, with only $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ being trainable. Despite the reduced number of parameters, as shown in the SFT performance in the table, performance remains stable even as $N_{B}$ increases. We attribute this to orthogonal and frozen $A$ and $B$, which maximize and preserve the expressiveness of PQ-LoRA (Theorem1), despite the lower trainable parameter count.
Method Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT ($N_{B}=2$) 66.14$\pm$0.02 57.02$\pm$0.09 47.75$\pm$0.42 46.02$\pm$0.14
SFT ($N_{B}=4$) 66.24$\pm$0.68 57.71$\pm$0.27 47.63$\pm$0.15 46.62$\pm$0.14
SFT ($N_{B}=8$) 66.21$\pm$0.62 57.87$\pm$0.20 47.86$\pm$0.29 46.76$\pm$0.20
FedMosaic ($N_{B}=2$) 68.07$\pm$0.10 59.90$\pm$0.68 51.07$\pm$0.06 48.78$\pm$0.15
FedMosaic ($N_{B}=4$) 67.86$\pm$0.51 59.83$\pm$0.16 51.26$\pm$0.04 49.36$\pm$0.08
FedMosaic ($N_{B}=8$) 67.52$\pm$0.77 59.39$\pm$0.23 51.27$\pm$0.27 49.26$\pm$0.14
A.17 Extended Fast Adaptation Evaluations
@@
A.18 Additional Details of DRAKE
We curate DRAKE using open-sourced datasets from the Internet. Specifically, we select large-scale multi-modal benchmarks that differ in distribution or properties from LLaVA’s pre-training data and provide clear metadata or instructions, enabling task-wise splits for emulating distribution shifts. We refer to diverse multi-modal benchmarks, such as DEMON (li2024demon, ), SEED-Bench-2 (li2024seed2, ), Co-Instruct (wu2024coinstruct, ), and HFLB (chen2024feddat, ).
@@ A.18 Additional Details of DRAKE
Visual relation tasks focus on capturing relationships among visual objects, consist of 12 distinct tasks. We categorize them into three splits: fashion-relation, dual-image relation, and spatial/temporal. The fashion-relation split focuses on fine-grained attribute understanding in clothing images, using the cloth-counting and three-/four-image variants of Fashion200K (han2017fashion200k, ) as well as the query–reference caption split of FashionIQ (wu2021fashioniq, ). The dual-image relation split evaluates comparative reasoning across separate images, drawing on NLVR2 (suhr-etal-2019nlvr, ), CIRR (liu2021cirr, ), the two- and four-image true/false subsets of VISION (bai2023vision, ), and MagicBrush (zhang2023magicbrush, ). The spatial/temporal split targets positional and temporal dependencies, combining single-frame spatial reasoning in VSR (liu2023vsr, ) with temporal understanding tasks from SEED-Bench-2 (li2024seed2, ).
@@
Multi-modal reasoning tasks require the integration of visual cues with common-sense knowledge, comprising 12 distinct tasks. We include four tasks from IRFL (yosef-etal-2023-irfl, ) that examine the figurative interpretation of images paired with non-literal language. We also include five additional tasks derived from the positive–negative image groups in Bongard-HOI (jiang2022bongardhoi, ) and Bongard-OpenWorld (wu2024bongardopenworld, ). These tasks require fine-grained discrimination between contrasting sets. We incorporate three reasoning challenges from COMICS (iyyer2017comics, ), MIT-States (isola2015mitstate, ), and VizWiz (gurari2018vizwiz, ), all sourced from DEMON (li2024demon, ).
@@
The VQA sub-group includes 16 diverse VQA tasks that differ significantly from LLaVA’s pre-training datasets. We include novel question types, such as image-quality queries in Co-Instruct (wu2024coinstruct, ), textbook-style diagram interpretation in TQA (kembhavi2017tqa, ), DVQA (kafle2018dvqa, ), and knowledge-grounded QA tasks from SEED-Bench-2 (li2024seed2, ). We also incorporate unfamiliar image domains, including IconQA (lu2021iconqa, ) and WCVQA (winata2024wcvqa, ). We partition Co-Instruct by question type (e.g., Yes/No, How/what questions), IconQA by difficulty and question category (e.g., Multi-choice or Short answering, kindergarten-level or grade 1-level questions), while we divide WCVQA using their original split.
Client 1 Client 2 Client 3 Client 4 Client 5
$\bullet$ Fashion200K (han2017fashion200k, ): ClothCombination
$\bullet$ FashionIQ (wu2021fashioniq, )
$\bullet$ VISION (bai2023vision, ): 2_Img
$\bullet$ MagicBrush (zhang2023magicbrush, )
$\bullet$ Co-Instruct (wu2024coinstruct, ): 2_ImgCompare
$\bullet$ SEED–Bench–2 (li2024seed2, ): KGQA
$\bullet$ IconQA (lu2021iconqa, ): ShortAnswerEasy
$\bullet$ WCVQA (winata2024wcvqa, ): DishName
$\bullet$ Co-Instruct (wu2024coinstruct, ): HowWhat
$\bullet$ DVQA (kafle2018dvqa, )
$\bullet$ IconQA (lu2021iconqa, ): ShortAnswerHard
$\bullet$ WCVQA (winata2024wcvqa, ): ContextDishName
$\bullet$ Co-Instruct (wu2024coinstruct, ): 3_ImgCompare
$\bullet$ SEED–Bench–2 (li2024seed2, ): InstanceQA
$\bullet$ IconQA (lu2021iconqa, ): MultiChoiceEasy
$\bullet$ WCVQA (winata2024wcvqa, ): AdvContextDishName
$\bullet$ Fashion200K (han2017fashion200k, ): ColorConsistency
$\bullet$ NLVR2 (suhr-etal-2019nlvr, ): Subset1
$\bullet$ VISION (bai2023vision, ): 4_Img
$\bullet$ VSR (liu2023vsr, )
Client 6 Client 7 Client 8 Client 9 Client 10
$\bullet$ Fashion200K (han2017fashion200k, ): StyleConsistency
$\bullet$ NLVR2 (suhr-etal-2019nlvr, ): Subset2
$\bullet$ CIRR (liu2021cirr, )
$\bullet$ SEED–Bench–2 (li2024seed2, ): TemporalQA
$\bullet$ IRFL (yosef-etal-2023-irfl, ): MetaphorSimileMatching
$\bullet$ Bongard–HOI (jiang2022bongardhoi, ): ActionDetection
$\bullet$ VizWiz (gurari2018vizwiz, )
$\bullet$ MIT–States (isola2015mitstate, )
$\bullet$ IRFL (yosef-etal-2023-irfl, ): FigurativeVerification
$\bullet$ IRFL (yosef-etal-2023-irfl, ): IdiomMatching
$\bullet$ Bongard–OpenWorld (wu2024bongardopenworld, ): ConceptDetection
$\bullet$ COMICS (iyyer2017comics, ): Dialogue
$\bullet$ IRFL (yosef-etal-2023-irfl, ): PhraseSelection
$\bullet$ Bongard–OpenWorld (wu2024bongardopenworld, ): ConceptQuery
$\bullet$ Bongard–HOI (jiang2022bongardhoi, ): ConceptQuery
$\bullet$ Bongard–HOI (jiang2022bongardhoi, ): ActionIncoherence
$\bullet$ Co-Instruct (wu2024coinstruct, ): YesNO
$\bullet$ TQA (kembhavi2017tqa, )
$\bullet$ IconQA (lu2021iconqa, ): MultiChoiceHard
$\bullet$ WCVQA (winata2024wcvqa, ): Location
@@
To assess generalization capability, we add 7 additional tasks that are disjoint from the clients’ tasks in DRAKE. We selected challenging benchmarks requiring extensive training, which are far from fast adaptation. The unseen tasks include novel task types (DreamSim (fu2023dreamsim, ), ImageCoDe (krojer-etal-2022-imagecode, )) and novel visual domains with familiar task format (RecipeQA (yagcioglu-etal-2018-recipeqa, ) subsets proposed in DEMON (li2024demon, ), and HQ-Edit hui2025hqedit ). In addition, we split the Contrast-Caption subset of Mantis (jiang2024mantis, ; yu2022coca, ) into two tasks: one with images overlapping LLaVA’s pre-training distribution and another with previously unseen images.
@@
A.19 Limitations and Future Work
While we evaluate FedMosaic across various multi-modal FL benchmarks, e.g., HFLB and our proposed DRAKE, and LLM benchmarks, e.g., Fed-Aya and Fed-LLM, computational constraints limit our experiments to a maximum of 10 clients. Exploring performance with a larger number of clients would better reflect real-world FL scenarios and represents an important direction for future work.
A.20 Impact Statements
@@
A.21 Detailed Algorithm of FedMosaic
@@
#     Number of clients $N$, Number of client model types $K$,
@@
#     Batch size $B$, Learning rate $\eta$, Gradient frequency $f$, EMA ratio $a$,
@@
#     Training data streams of $T$ tasks for $N$ clients $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{N}\}$,
@@
                               g_t.append(LastLayerGrad($\mathcal{W}_{S}$, loss)
@@
      ## Step 2. Alignment of A using MSE loss
@@
            # Compute MSE loss for each layer using outputs captured by hooks
@@
                   # Compute MSE between LoRA A outputs of W_i and W_p at layer $j$
                   loss_mse = MSE(W_p.lora_outputs[j], W_i.lora_outputs[j])
@@
                   total_loss += loss_mse + $\lambda$ loss_reg
[…]
      ## Step 3. Post-process A to enforce Orthogonality
@@
       ## Step 1. Attach hooks to LoRA B and perform a forward pass
@@
             # Compute CCA projection vectors
@@
             # Initialize B using the CCA transformation
             W_i[j].B = (($\Pi$_i)-1)T · ($\Pi$_p)T · W_p[j].B

### v3

@@ A B $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
Table 32: Accuracy on different weight alignment in PQ-LoRA under Heterogeneous PFL-Static setup on DRAKE. ‘Random’ initializes matrices randomly, while Aligned (L2) and Aligned (CCA) use matrices aligned by L2 loss and canonical correlation analysis (CCA), respectively.
A.22 Comparison with Similarity-Aware Model Aggregation
Federated Learning
We are not the first to try to aggregate client models based on task similarity, but most existing approaches are not applicable to large language model (LLM) federated learning scenarios. pFedSim (Tan et al., 2023) measures the classifier similarity among clients and uses it for weighted model aggregation, which is infeasible for LLM fine-tuning where the pre-trained classifier remains frozen. pFedHR (Wang et al., 2023) computes similarity between client models based on the output logits and performs similarity-aware layer stitching. However, its logit extractions relies on public data, which (as acknowledged by the authors) raises public data sensitivity concerns: the discrepancy between private and public data may leads to inaccurate similarity estimate. Flashback (Aljahdali et al., 2024) conducts dynamic knowledge distillation where teacher logits are adaptively weighted by client-wise label count, but this method is limited to single-dataset classification FL tasks. There is also another approach like FEDLAW (Li et al., 2023d) that adaptively aggregates models based on the optimized learnable weights instead of client similarity. However, it incurs high computational costs and remains highly sensitive to the choice of public data.
RELA, on the contrary, measures the task similarity among clients using sanitized last layer gradients. These gradients from each private data effectively capture task characteristics with minimal additional computation and communication cost, while preserving privacy through EMA updates, noise injection, and gradient compression.
Multi-Task Learning
Federated learning (FL) and personalized federated learning (PFL) are not the only paradigms that learn heterogeneous tasks concurrently. Multi-task learning (MTL) trains a single model on multiple tasks simultaneously in a single compute node, not in separate private nodes as in FL. MTL methods, such as NBS (Navon et al., 2022) and Rotograd (Javaloy & Valera, 2022), leverage inter-task relationships to dynamically weight gradients, thereby mitigating interference and promoting a balanced update direction across tasks. This is similar to the motivation of our aggregation method, i.e., RELA, which aims to reduce the interference between multiple tasks. However, these techniques are not directly applicable to personalized federated learning due to differing objectives: similarity-aware MTL strategies target generalizability by reshaping task-specific gradients to improve a single shared model, whereas RELA focuses on personalization. It provides each client with a customized global model by down-weighting contributions from irrelevant tasks, rather than enforcing a single globally shared one, thereby providing knowledge beneficial for personalization, while still enhancing generalizability as a byproduct.
We also empirically compare RELA with MTL methods (Navon et al., 2022). Note that Rotograd(Javaloy & Valera, 2022) is incompatible with PFL because it aligns gradient directions by optimizing rotation heads using samples from multiple tasks, which is not feasible under federated data isolation and privacy constraints. Therefore, we only compare with NBS, which determines task weights via a Nash bargaining objective that maximizes total loss reduction. In PFL setting, we apply NBS weighting to adapter parameters rather than raw gradients for a shared global model, since FL aggregates parameters. As shown in Tab. 33, NBS shows lower ’Self’ performance, while maintaining similar generalization (‘Others’) performance. We believe NBS’s weighting mechanism focuses on harmonizing parameters from multiple tasks to minimize the overall loss, but fails to preserve or share task-relevant information necessary for each client’s personal tasks, resulting in poor ‘Self’ performance.
Self Others
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 33: Quantitative comparison with MTL method.
A.23 Extended Fast Adaptation Evaluations
@@
A.24 Additional Details of DRAKE
We curate DRAKE using open-sourced datasets from the Internet. Specifically, we select large-scale multi-modal benchmarks that differ in distribution or properties from LLaVA’s pre-training data and provide clear metadata or instructions, enabling task-wise splits for emulating distribution shifts. We refer to diverse multi-modal benchmarks, such as DEMON (Li et al., 2024c), SEED-Bench-2 (Li et al., 2024a), Co-Instruct (Wu et al., 2024b), and HFLB (Chen et al., 2024).
@@ A.24 Additional Details of DRAKE
Visual relation tasks focus on capturing relationships among visual objects, consist of 12 distinct tasks. We categorize them into three splits: fashion-relation, dual-image relation, and spatial/temporal. The fashion-relation split focuses on fine-grained attribute understanding in clothing images, using the cloth-counting and three-/four-image variants of Fashion200K (Han et al., 2017) as well as the query–reference caption split of FashionIQ (Wu et al., 2021). The dual-image relation split evaluates comparative reasoning across separate images, drawing on NLVR2 (Suhr et al., 2019), CIRR (Liu et al., 2021), the two- and four-image true/false subsets of VISION (Bai et al., 2023), and MagicBrush (Zhang et al., 2023a). The spatial/temporal split targets positional and temporal dependencies, combining single-frame spatial reasoning in VSR (Liu et al., 2023a) with temporal understanding tasks from SEED-Bench-2 (Li et al., 2024a).
@@
Multi-modal reasoning tasks require the integration of visual cues with common-sense knowledge, comprising 12 distinct tasks. We include four tasks from IRFL (Yosef et al., 2023) that examine the figurative interpretation of images paired with non-literal language. We also include five additional tasks derived from the positive–negative image groups in Bongard-HOI (Jiang et al., 2022) and Bongard-OpenWorld (Wu et al., 2024d). These tasks require fine-grained discrimination between contrasting sets. We incorporate three reasoning challenges from COMICS (Iyyer et al., 2017), MIT-States (Isola et al., 2015), and VizWiz (Gurari et al., 2018), all sourced from DEMON (Li et al., 2024c).
@@
The VQA sub-group includes 16 diverse VQA tasks that differ significantly from LLaVA’s pre-training datasets. We include novel question types, such as image-quality queries in Co-Instruct (Wu et al., 2024b), textbook-style diagram interpretation in TQA (Kembhavi et al., 2017), DVQA (Kafle et al., 2018), and knowledge-grounded QA tasks from SEED-Bench-2 (Li et al., 2024a). We also incorporate unfamiliar image domains, including IconQA (Lu et al., 2021) and WCVQA (Winata et al., 2024). We partition Co-Instruct by question type (e.g., Yes/No, How/what questions), IconQA by difficulty and question category (e.g., Multi-choice or Short answering, kindergarten-level or grade 1-level questions), while we divide WCVQA using their original split.
Client 1 Client 2
$\bullet$ Fashion200K (Han et al., 2017): ClothCombination
$\bullet$ FashionIQ (Wu et al., 2021)
$\bullet$ VISION (Bai et al., 2023): 2_Img
$\bullet$ MagicBrush (Zhang et al., 2023a)
$\bullet$ Co-Instruct (Wu et al., 2024b): 2_ImgCompare
$\bullet$ SEED–Bench–2 (Li et al., 2024a): KGQA
$\bullet$ IconQA (Lu et al., 2021): ShortAnswerEasy
$\bullet$ WCVQA (Winata et al., 2024): DishName
Client 3 Client 4
$\bullet$ Co-Instruct (Wu et al., 2024b): HowWhat
$\bullet$ DVQA (Kafle et al., 2018)
$\bullet$ IconQA (Lu et al., 2021): ShortAnswerHard
$\bullet$ WCVQA (Winata et al., 2024): ContextDishName
$\bullet$ Co-Instruct (Wu et al., 2024b): 3_ImgCompare
$\bullet$ SEED–Bench–2 (Li et al., 2024a): InstanceQA
$\bullet$ IconQA (Lu et al., 2021): MultiChoiceEasy
$\bullet$ WCVQA (Winata et al., 2024): AdvContextDishName
Client 5 Client 6
$\bullet$ Fashion200K (Han et al., 2017): ColorConsistency
$\bullet$ NLVR2 (Suhr et al., 2019): Subset1
$\bullet$ VISION (Bai et al., 2023): 4_Img
$\bullet$ VSR (Liu et al., 2023a)
$\bullet$ Fashion200K (Han et al., 2017): StyleConsistency
$\bullet$ NLVR2 (Suhr et al., 2019): Subset2
$\bullet$ CIRR (Liu et al., 2021)
$\bullet$ SEED–Bench–2 (Li et al., 2024a): TemporalQA
Client 7 Client 8
$\bullet$ IRFL (Yosef et al., 2023): MetaphorSimileMatching
$\bullet$ Bongard–HOI (Jiang et al., 2022): ActionDetection
$\bullet$ VizWiz (Gurari et al., 2018)
$\bullet$ MIT–States (Isola et al., 2015)
$\bullet$ IRFL (Yosef et al., 2023): FigurativeVerification
$\bullet$ IRFL (Yosef et al., 2023): IdiomMatching
$\bullet$ Bongard–OpenWorld (Wu et al., 2024d): ConceptDetection
$\bullet$ COMICS (Iyyer et al., 2017): Dialogue
Client 9 Client 10
$\bullet$ IRFL (Yosef et al., 2023): PhraseSelection
$\bullet$ Bongard–OpenWorld (Wu et al., 2024d): ConceptQuery
$\bullet$ Bongard–HOI (Jiang et al., 2022): ConceptQuery
$\bullet$ Bongard–HOI (Jiang et al., 2022): ActionIncoherence
$\bullet$ Co-Instruct (Wu et al., 2024b): YesNO
$\bullet$ TQA (Kembhavi et al., 2017)
$\bullet$ IconQA (Lu et al., 2021): MultiChoiceHard
$\bullet$ WCVQA (Winata et al., 2024): Location
@@
To assess generalization capability, we add 7 additional tasks that are disjoint from the clients’ tasks in DRAKE. We selected challenging benchmarks requiring extensive training, which are far from fast adaptation. The unseen tasks include novel task types (DreamSim (Fu et al., 2023b), ImageCoDe (Krojer et al., 2022)) and novel visual domains with familiar task format (RecipeQA (Yagcioglu et al., 2018) subsets proposed in DEMON (Li et al., 2024c), and HQ-Edit Hui et al. (2025)). In addition, we split the Contrast-Caption subset of Mantis (Jiang et al., 2024; Yu et al., 2022) into two tasks: one with images overlapping LLaVA’s pre-training distribution and another with previously unseen images.
@@
A.25 Comparison with Other FL Benchmarks.
A number of federated learning (FL) benchmarks exist, but most lack critical aspects required for FL of multi-modal foundation models, which DRAKE addresses, with the main differences summarized in Tab. 1.
Benchmarks like NonIID-50 (Yoon et al., 2021), LEAF-FCL (Qi et al., 2023), and MNIST-Shuffle (Wuerkaixi et al., 2024) simulate distribution shifts under FL. However, they are single-task and single-modal (i.e., image classification) only, thus cannot reflect real-world task diversity. HC-FMTL (Lu et al., 2024) spans multiple vision tasks, such as depth estimation and semantic segmentation, but still remains unimodal.
On the language side, several text-only FL benchmarks have recently been proposed to target LLM federated learning scenarios. Fed-SNI (Collins et al., 2023) and Fed-FLAN (Long et al., 2024) cover diverse NLP tasks, while Fed-Aya (Ye et al., 2024) focuses on multilingual instruction following. FEDLEGAL (Zhang et al., 2023b) curates a FL benchmark for the privacy-sensitive legal domain where federated learning is necessary. Despite their breadth, they are also limited to a single modality and assume static data distribution, while data distribution often shifts over time in real-world.
For multi-modal PFL, we find only one prior benchmark: HFLB (Chen et al., 2024). Although HFLB includes 7 different datasets, the multi-modal task diversity is limited, and they are mostly single-image settings. Other multi-modal FL benchmarks, e.g., FedMultimodal (Feng et al., 2023), FHBench (Wang et al., 2025), and FedMLLM (Xu et al., 2024), are primarily comprised of only a classification task on different modalities or split one or two VQA datasets, which is insufficient for evaluation personalization under diverse heterogeneous tasks. Importantly, existing benchmarks generally lack an unseen task split to test the generalization ability of the model.
In contrast, DRAKE offers 40 different multi-modal tasks (including multi-image tasks), which is suitable for personalization evaluation and temporal distribution shift. Moreover, DRAKE additionally provides 7 unseen tasks for out-of-distribution generalization evaluation.
A.26 Limitations and Future Work
While we evaluate FedMosaic across diverse multi-modal FL benchmarks (e.g., HFLB and our proposed DRAKE) and text-only NLP benchmarks (e.g., Fed-Aya, Fed-Scope, and Fed-LLM-Large), it has not yet been tested on other modalities such as speech and time series. Extending evaluations to these domains would better capture real-world FL scenarios and represent an important direction for future work.
Moreover, although LoRA is most commonly used in transformer-based models, it is not restricted to them. Specifically, prior studies (Yeh et al., 2023; Ding et al., 2024) have demonstrated its applicability to convolutional layers, highlighting its architecture-agnostic nature. Therefore, exploring the use of PQ-LoRA beyond transformers to enable knowledge transfer across broader model families represents an important direction for future work.
A.27 Impact Statements
@@
A.28 Detailed Algorithm of FedMosaic
@@
#   Number of clients NN, Number of client model types $K$,
@@
#   Batch size $B$, Learning rate $\eta$, Gradient frequency $f$, EMA ratio $a$, Temperature $\tau$, Gaussian noise std $\mu$, Gradient subsample ratio $N_{s}$
@@
#   Training data streams of $T$ tasks for NN clients $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{N}\}$,
@@
1ex
# Set Random subsample indices for last layer gradient
subsample_index = random_choice(last_layer_grad size, $N_{s}$)
1ex
@@
              g_last = LastLayerGrad($\mathcal{W}_{S}$, loss)
              # Subsample gradient and add gaussian noise
              eps $\sim$ $N(0,I)$ # $I\in\mathbb{R}^{\text{last\_layer\_grad size}}$
              g_sanitized = g_last[subsample_index] + $\mu$ * eps[subsample_index]
              g_t.append(g_sanitized)
@@
 1ex## Step 2. Alignment of A using L2 loss
@@
    # Compute L2 loss for each layer using outputs captured by hooks
@@
        # Compute L2 between LoRA A outputs of W_i and W_p at layer $j$
        loss_l2 = L2(W_p.lora_outputs[j], W_i.lora_outputs[j])
@@
        total_loss += loss_l2 + $\lambda$ * loss_reg
    W_i -$\mathrel{-}=$ $\eta$ * Grad(W_i, total_loss)
 1ex## Step 3. Post-process A to enforce Orthogonality
@@
 1ex## Step 1. Attach hooks to LoRA B and perform a forward pass
@@
    1ex# Compute CCA projection vectors
@@
    1ex# Initialize B using the CCA transformation
     W_i[j].B = $($$\Pi$_i)-1)T · ($\Pi$_p)T · W_p[j].B
1ex